OpenStudy (anonymous):

please helppp!! the graph of which function passes throught the points (-2, 2)?? thanks :)

OpenStudy (anonymous):

you have no idea how many functions can pass through this point :P I'm sure you have choices right?

OpenStudy (anonymous):

a. y= -x^2 + 6 b. y= -x^2 -6 c. y= x^2 + 6 d. y= 6x^2 -1 lol sorry for not posting these before, thanks for your help tho

OpenStudy (anonymous):

lol it's okay ,to be honest the question is solved by substituting x=-2 and y=2 and see which one will has equal sides

OpenStudy (anonymous):

not that hard right :P ?

OpenStudy (anonymous):

thank you soo much!!! :)

OpenStudy (anonymous):

you are welcome so much :)